OpenStudy (anonymous):

A 10 kg box is being pulled by a force F = 85 N at an angle of 57degrees from the horizontal. Find the following quantities if the box is accelerated at 2.5 m/s2 for 14 m? a. Work done by the external force b. Work done by frictional force c. Work done by net force

OpenStudy (anonymous):



OpenStudy (anonymous):

Work Formula: \[Work = Force \times Distance\]

OpenStudy (anonymous):

Hi! I understand how to do a I just don't understand b and c

OpenStudy (anonymous):

Any help would be great!

OpenStudy (abmon98):

Workdone by external force=Force*Distance W=85*14=1190 Frictional Force is the Force Opposite to the force of movement R is the component of a force at right angles to the surface R=10gcos(57) External Force-(FrictionalForce+mgsin*angle)=m(mass)*a(acceleration) 85-(Ff+(10(g)*sin57)=10*2.5 85-(100*sin57)-ff=25

OpenStudy (anonymous):

Okay wow, thanks, how about the work done by the net force?

OpenStudy (abmon98):

Your most welcome I think that net force is calculated using (external force-frictional force)